OpenStudy (anonymous):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

OpenStudy (anonymous):

Do you know the formula for equation of a line?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

First thing is you have to find the slope

OpenStudy (anonymous):

slope = second y value from the first y value / divided by the second x value from the first x value

OpenStudy (anonymous):

once you have the slope, you can use plug that into the formula: y-y1=m(slope)(x-x1)

OpenStudy (anonymous):

y = -2x + 1 y = 2x + 1 correct? y = 1/2 x - 1 y = -1/2 x + 1

OpenStudy (anonymous):

@jtryon